OpenStudy (anonymous):

A can of soup is 3 inches in diameter and 5 inches in height. How much paper is needed to make a label for the soup can?

OpenStudy (anonymous):

\[2\pi(\frac{ 3 }{ 2})(5)=?\]

OpenStudy (anonymous):

this comes from \[2\pi r h\] which will give you the label around the soup can